Question 139 of 417MediumMCQ
Let N denote the sum of numbers obtained when two dice are rolled. If the probability that 2N<N! is nm (coprime), find 4m−3n. Solutions · 1
Official solution
2N<N! holds for N≥4.
P(N=2)=361, P(N=3)=362, so P(N<4)=363.
\[P(N\ge4) = 1-\frac{3}{36}=\frac{33}{36}=\frac{11}{12} \implies m=11,n=12\]
4m−3n=44−36=8 Answer the question to unlock